# Chiral nature of current flow in the regime of the quantized Hall effect

### v2

@@ III.1 Sample and measurement setup
Figure 5: Three-terminal measurement with symmetric circuit, cf. inset in (a). A voltage of $V=1\,$mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating. $\vec{B}$ points from the surface into the 2DES plain such that $I_{xy}$ is bent towards contacts 2 and 3. (a) Current ratios $-I_{3}/I_{1}$ and $-I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq. (II.2) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq. (II.2) with $\mu(B)$ as free parameter (dotted red lines). (b) enlargements near $\nu=1$ showing experimental data, the Drude model prediction for $\mu=\mu_{0}$ and the model prediction of Eq. (4) for coherent and chiral current at $\nu=1$ (dotted red lines).
@@
For multi-terminal current measurements, we apply the voltage of $V_{1}=\pm 1\,$mV to the source contact 1 (using a Keithley 2450 source meter). We measure the current $I_{1}(B)$ flowing from the Hall bar into contact 1, using the source meter, and the currents $I_{3}(B)$ and $I_{5}(B)$ flowing into contacts 3 or 5, respectively, using Basel Precision Instr. current amplifiers. Contacts 3 and 5 are connected to the electrical ground, while the remaining contacts 2, 4, and 6 may be either electrically floating or connected to ground depending on the experiment, cf. Fig. 4. Here, we apply $B$ perpendicular to the 2DES, such that it points from the sample surface towards the 2DES, hence, electrons contributing to $I_{xy}$ move from contact to contact along the counterclockwise pathway $1\to 2\to 3\to 4\to 5\to 6\to 1$. Due to the direction of the magnetic field and a voltage applied to contact 1, the largest portion of $I_{xy}$ will always flow between contact 1 and contact 3, if contact 2 is floating, or to contact 2, if it is grounded. Whether this current is positive or negative depends on the sign of $V_{1}$.
We performed a series of measurements at $B=0$ to determine the contact resistances ($R_{i}\sim 2.5\,\text{k}\Omega$) and offset voltages ($V_{3,5}\sim 20\,\mu$V) of the two current amplifiers. In addition, we used the comparison between measured data and model curves for a fine tuning of these circuit parameters, which remain constant between measurements. They are listed in Table 1.
contact $i$ resistance $R_{i}$ ($\Omega$) voltage $V_{i}$ (mV) geometry $a_{i}$
1 2420 1.000 1
2 2420 0 0.5
3 2400 -0.021 1
4 2350 0 1
5 2420 -0.018 1
6 2400 0 0.5
Table 1: Parameters of the circuit sketched in Fig. 4. Contact resistances $R_{i}$ include ohmic contacts, cables and $RC$-filters. Voltage $V_{1}$ is applied to contact 1, $V_{2}$ and $V_{3}$ are offset voltages of the current amplifiers. To account for the contribution of the Hall bar itself to the resistance, we defined a geometry parameter and use $R_{i}\to R_{i}+a_{i}/[n_{\text{s}}e\mu(B)]$ in the model calculations of the ohmic contributions. The correction is negligible for $B=0$.
III.2 Three-terminal measurements
@@
First, we discuss a three-terminal measurement in a mirror symmetric geometry, for which we keep the contacts 2, 4 and 6 electrically floating; current flows from contact 1 into contacts 3 and 5. In Fig. 5(a), we present as solid (blue) lines the measured current ratios $-I_{3}(B)/I_{1}(B)$ and $-I_{5}(B)/I_{1}(B)$. As expected, we find that $|I_{3}|$ increases with $B$ while $|I_{5}|$ decreases accordingly, such that $I_{3}(B)+I_{5}(B)=-I_{1}(B)$. The dashed black line indicates the prediction of the model according to Eq. (II.2) for $n=3$ if we assume a constant mobility $\mu(B)=\mu_{0}\equiv\mu(B=0)=395\,\textrm{m}^{2}/\textrm{Vs}$. For this large mobility, the Ohmic term (first term) of Eq. (II.2) can be neglected for $B\gtrsim 50\,$mT (corresponding to $\mu B\gtrsim 20$), such that with $(\nu G_{0})^{-1}=R_{\text{H}}(B)$ Eq. (4) [Drude model] and Eq. (II.2) [chiral model] become identical.
@@
In the classical limit for $B\lesssim 0.5$ T the Drude model agrees well with our measurement. As we increase $B$ further, the Drude model still predicts the correct trend but as long as we use a constant $\mu(B)=\mu_{0}$ (black dashed lines), we observe growing deviations. These are clearly visible in the enlargement around the first plateau at $\nu=1$ in Fig. 5(b). The red dashed lines are the prediction of the chiral model according to Eq. (4) [or Eq. (II.2) using $\mu B=\infty$], where we inserted a constant filling factor $\nu=1$. It perfectly fits the plateau values of the current ratios. An exact match between measurements and model beyond the plateaus can be achieved by using $\mu(B)$ as a free fit parameter in Eq. (II.2). For the dotted red line in Fig. 5(a), we used the magnetic field dependent values of $\mu(B)$ as shown by red squares in Fig. 6. For comparison, we have also determined $\mu(B)$ from the Shubnikov-de-Hass oscillations of a standard longitudinal resistance measurement, see the black dots in Fig. 6.  [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to $T>4\,$K between the two measurements. The general agreement between the $\mu(B)$ values obtained from longitudinal resistance measurements with those obtained by fitting Eq. (II.2) to our three-terminal current measurements supports the validity of our model.
@@
The two regions of divergent $\mu(B)$ in Fig. 6 correspond to the plateau regions with the local filling factors $\nu=2$ and $\nu=1$. Here, $I_{xx}$ vanishes. Away from the plateau regions, we find strongly reduced mobility values, indicating a transition to diffusive transport. The difference between the Drude model predictions for $\mu=\mu_{0}$ [dashed line in Fig. 5(a)] and the measured currents is the ohmic current contribution $I_{xx}$, which becomes clearly visible for $\mu<20$.
@@
So-far we have discussed three-terminal current measurements. With the knowledge of the contact resistances of our sample, we can quantitatively compare the experiments with model predictions. Between the plateaus, the mobility is reduced and the Drude model fits perfectly to our data, indicating non-chiral transport. Regarding the plateau regions, our three-terminal measurements are clearly consistent with chiral transport. Importantly, any additional non-zero transmission coefficients in Eq. (3) due to non-chiral currents would modify the prediction of Eq. (4). Clearly, we do not observe such deviations.
@@ III.3 Four-terminal measurements
In Fig. 7, we present current ratios $-I_{i}/I_{1}$ for three different four-terminal measurements (blue solid lines). Again, we measure the currents into contacts $i=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig. 7(a), contact 6 in (b), and contact 2 in (c). If we assume $\mu(B)=\mu_{0}$, Eq. (II.2) reproduces the overall dependences $I_{i}(B)$ with a similar accuracy and similar deviations as it did for the three-terminal measurement (dashed black lines). The dotted red lines correspond to the predictions of Eq. (II.2), where we used the identical $\mu(B)$ that we found by fitting the three-terminal measurement above, cf. Fig. 6. As it was the case for the three terminal measurements, for the plateau regions, corresponding to diverging $\mu(B)$, the agreement between theory and the measured currents is perfect. However, away from the plateaus, the theory curves using the varying $\mu(B)$ does no longer yield a perfect agreement between theory and measured data. It means, that the division of the ohmic $I_{xx}$ component of the current is not accurately predicted.
Nevertheless, our four-terminal measurements confirm the main conclusions of our three-terminal measurements: (1) Away from the plateaus, the transport is not chiral. (2) For the plateau regions, the agreement between our four-terminal measurements and the chiral and coherent model supports chiral transport.
@@ III.4 Limits of our model in the diffusive regime
Away from the plateaus, the measured currents depend on the contact configuration beyond the predictions of our model. We conjecture, that our initial assumption of a homogeneous Hall bar fails near the intersections. In the regime of diffusive transport it implies a homogeneous electric field inside the Hall bar. In reality, near intersections, both components, $E_{x}$ and $E_{y}$ of the electric field and, therefore, the currents $I_{xx}$ and $I_{xy}$ become inhomogeneous. While $I_{xy}$ is still completely bent into the contacts, the division of $I_{xx}$ is influenced by its inhomogeneity. A more realistic model would require a suitable numerical calculation, e.g., based on a path integral formalism. We conjecture, that such a corrected model would yield perfect agreement with measurements, namely that the Drude model is still correct for more than three terminals, where it worked perfectly for three terminals.
III.5 Purely experimental confirmation of non-chiral currents between plateaus
@@
Figure 8: Comparison of the respective largest measured current contributions flowing into a grounded contact for three different configurations of four-terminal measurements. The dotted horizontal lines depict the predicted current ratios for filling factors $1\leq\nu\leq 4$. The current ratios $I_{i}\,/\,I_{1}$ marked with a * are slightly scaled to account for the slightly different contact resistances, such that all three theory curves according to Eq. (II.2) with $\mu=\mu_{0}$ are identical and follow the edge between shaded and white backgrounds. Sketches: Arrows indicate the displayed current branches; colored numbers indicate current carrying and gray numbers floated contacts.
we plot for each of the three four-terminal measurements shown in Fig. 7 only the branch $I_{k}/I_{1}$ corresponding to the largest current. In a chiral model, this is the current flowing into the first grounded contact after emission from contact 1. We slightly scaled the data (based on the values provided in Table 1) to compensate for small variations in the contact resistances. After scaling, the three curves are identical in the classical regime ($B<0.5\,$T) as well as on the plateaus of the QHE. The line linking shaded and white backgrounds corresponds to the prediction of Eq. (II.2) for constant $\mu=\mu_{0}$ [equal to the dashed black lines in Fig. 7(a)].
Compared to the model prediction, between the plateaus the current is reduced for all three geometries, which above we already interpreted in terms of a finite contribution of the non-chiral $I_{xx}$. Because the various contact resistances of our sample are almost identical, for a purely chiral current flow the branches of the largest current $I_{k}(B)/I_{1}$, shown in Fig. 8 should be independent of the details of the contacts configuration (as long as the overall number of grounded contacts remains the same). The fact that $I_{k}(B)/I_{1}$ varies between configurations, is a direct experimental prove of non-chiral currents between the quantized plateaus.
This conclusion experts additional support by a closer look at the amounts of current reduction: It systematically depends on the relative distances between contact 1 and the grounded contacts on both sample edges in a way that contradicts chiral current flow.
@@ III.6 Formulation of the proof of chiral current flow for quantized Hall resistance
So far, we have proven non-chiral current flow between the quantized plateaus, while for the quantized plateaus we have merely stated that our experimental results are consistent with chiral current flow. However, our analysis allows a stronger statement, namely that the agreement between the chiral model and our measurements is a prove of chiral current flow.
The proof is based on the coherent and chiral model introduced in Sec. II.1. Based on the Landauer-Büttiker formalism it describes the current division into $n$ contacts in terms of the mutual conductance coefficients $G_{ij}$ between the contacts. In this model, chiral current flow corresponds to the specific values of conductance coefficients given by Eq. (3). A non-chiral behavior would correspond to different values, including additional non-zero conductance coefficients in Eq. (3). These would inevitably result in different plateau values of the predicted current ratios.
In all our measurements including three or four terminals shown in Fig. 5, Fig. 7 and Fig. 8, we instead see a perfect agreement between our chiral model prediction and the measured plateau values of the current ratios. This is in particular evident when comparing the horizontal dotted lines in Fig. 5(b) and Fig. 8, which indicate the predictions of Eq. (4) for the corresponding integer filling factors, with the corresponding measured plateau values.
This result stands in contrast to our finding in between plateaus, where the deviations of our measured currents from the prediction of Eq. (II.2) for $\mu B\gg 1$ provide a proof of non-chiral currents.
@@ IV Summary and Outlook
We have performed multiterminal current measurements of a Hall bar in the regime of the QHE. We find that the current flow is chiral in the magnetic field regions of the quantized Hall resistance plateaus but becomes non-chiral in between the plateaus. Our result is based on the direct comparison of the measured current distribution into various contacts (three or four terminals) with two limiting models, which offer well understood, unique physical meanings. The classical Drude model serves us as the basis to describe non-chiral diffusive transport, while our second model is based on the generic chiral and coherent transmission problem. The comparison of different contact configurations for more than three terminals allows for a direct experimental proof of non-chiral currents, which confirm our model based results. Our work contributes to the understanding of the current flow inside mesoscopic devices in the regime of the QHE. It thereby confirms the assumption of chiral current flow frequently used for the case of quantized plateaus of the Hall resistance.

### v3

@@ III.2 Three-terminal measurements
Figure 7: Three-terminal measurement with symmetric circuit, cf. inset in (a). A voltage of V=1V=1\,mV is applied to contact 1, while contacts 3 and 5 are connected to ground via current amplifiers; contacts 2, 4, and 6 are electrically floating. $\vec{B}$ points from the surface into the 2DES plain such that $I_{xy}$ is bent towards contacts 2 and 3. (a) Current ratios $I_{3}/I_{1}$ and $I_{5}/I_{1}$ as a function of magnetic field $B$ (solid blue lines); Drude model prediction of Eq. (II.2) assuming constant $\mu=\mu_{0}$ (dashed black line); fit of Drude model prediction of Eq. (II.2) with $\mu(B)$ as free parameter (dotted red lines). (b) enlargements near $\nu=1$ showing experimental data, the Drude model prediction for $\mu=\mu_{0}$ and the model prediction of Eq. (4) for coherent and chiral current at $\nu=1$ (dotted red lines).
@@
we present as solid (blue) lines the measured current ratios $I_{3}(B)/I_{1}(B)$ and $I_{5}(B)/I_{1}(B)$. As already discussed above, $I_{3}/I_{1}$ increases with $B$ while $I_{5}/I_{1}$ decreases accordingly, such that $I_{3}(B)/I_{1}(B)-I_{5}(B)/I_{1}(B)=1$. The dashed black line indicates the prediction of the Drude model according to Eq. (II.2) for $n=3$, if we use the constant mobility $\mu(B)=\mu_{0}=395\,\textrm{m}^{2}/\textrm{Vs}$ that we measured at $B=0$.
@@
In the classical limit for $B\lesssim 0.5$ T the Drude model agrees well with our measurement. As we increase $B$ further, the Drude model still predicts the correct trend but as long as we use a constant $\mu(B)=\mu_{0}$ (black dashed lines), we observe growing deviations. These are clearly visible in the enlargement around the first plateau at $\nu=1$ in Fig. 7(b). The red dotted lines in Fig. 7(b) are the prediction of the chiral model according to Eq. (4) [corresponding to Eq. (II.2) for $\mu B=\infty$] and a constant filling factor $\nu=1$. It perfectly fits the values of the current ratios along the $\nu=1$ plateau of the quantized Hall resistance. An exact match between measurements and model beyond the plateaus can be achieved by using $\mu(B)$ as a free fit parameter in Eq. (II.2). For the dotted red line in Fig. 7(a), we used the magnetic field dependent values of $\mu(B)$, which are plotted as red squares in Fig. 8.
For comparison, we have also determined $\mu(B)$ from the SdH oscillations of a standard longitudinal resistance measurement, see the black dots in Fig. 8  [^1] The slight deviations between the two are likely related with local changes in the sample during thermal cycles of the He-3 evaporation cryostat to T>4T>4\,K between the two measurements.. The general agreement between the $\mu(B)$ values obtained from two completely independent measurements, namely longitudinal resistance measurements versus fitting Eq. (II.2) to our three-terminal current measurements, strongly supports the validity of our model.
@@
The two regions of divergent $\mu(B)$ in Fig. 8 correspond to the plateau regions with the local filling factors $\nu=2$ and $\nu=1$. Here, $I_{xx}$ vanishes supporting $\mu B\to\infty$. Away from the plateau regions, in Fig. 8 we find mobility values strongly reduced even compared to $\mu_{0}$. This indicates a transition from ballistic transport on the plateaus to diffusive transport between the plateaus. The difference between the Drude model predictions for $\mu=\mu_{0}$ [dashed line in Fig. 7(a)] and the measured currents [blue line] is precisely the ohmic current contribution $I_{xx}$, which becomes clearly detectable for $\mu B<20$.
To understand the meaning of the perfect agreement of the Drude model fits based on a variable $\mu(B)$ it is useful to mention, that for a large enough mobility, here $\mu B\gtrsim 20$, the Ohmic term (first term) of Eq. (II.2) becomes too small to be detected. Neglecting $I_{xx}$ in Eq. (II.2) and using $(\nu G_{0})^{-1}=R_{\text{H}}(B)$, then makes Eq. (4) [chiral model] and Eq. (II.2) [Drude model] mathematically identical.
@@
So-far we have discussed three-terminal current measurements. With the knowledge of the contact resistances of our sample, we can quantitatively compare the experiments with model predictions. Between the plateaus, the mobility is strongly reduced and the Drude model fits perfectly to our data, indicating non-chiral diffusive transport. Regarding the quantized plateau regions, our three-terminal measurements are clearly consistent with Eq. (4), describing chiral current flow, if we assume a constant integer filling factor (for the entire plateau). It confirms, that for the quantized Hall plateaus the current flows within regions of constant integer filling factor, while the bulk filling factor (averaged over the width of the Hall bar) continuously decreases proportional to $1/B$. Importantly, any additional non-zero transmission coefficients in Eq. (3) due to non-chiral currents would modify the prediction of Eq. (4). Clearly, the fact that Eq. (4) perfectly fits our data for the plateaus with $\mu B\to\infty$ (without the need of additional non-zero transmission coefficients) suggests chiral transport for the quantized Hall plateaus.
@@ III.3 Four-terminal measurements
Three-terminal measurements are special, because if a model curve agrees with one current branch, say $I_{3}/I_{1}$ in Fig. 7, because of $I_{1}+I_{3}+I_{5}=0$ it will also agree with the second branch, here $I_{5}/I_{1}$. Clearly, this is not the case if we add additional terminals. Then, any inaccuracies of our model will be visible as deviations between measurements and model predictions. In that sense, experiments with four or more terminals provide a stricter test of our models compared to the three-terminal measurements above.
In Fig. 9, we present current ratios $I_{k}/I_{1}$ for three different four-terminal measurements (blue solid lines). Again, we measure the currents into contacts $k=1,3$ and 5, but now we ground an additional contact, namely contact 4 in Fig. 9(a), contact 6 in (b), and contact 2 in (c). If we assume $\mu(B)=\mu_{0}$, Eq. (II.2) reproduces the overall dependences $I_{k}(B)$ with a similar accuracy and similar deviations as it did for the three-terminal measurement (dashed black lines).
The dotted red lines correspond to model curves of Eq. (II.2) using the $\mu(B)$ plotted in Fig. 8 and determined by fitting the three-terminal measurement shown in Fig. 7. As it was the case for the three terminal measurements, for the plateau regions, corresponding to diverging $\mu(B)$, the agreement between theory and the measured currents is perfect. This observation supports our previous conclusion, that in the plateau regions the current flow is chiral, while away from the plateaus, the chiral model fails.
Noticeable, away from the plateaus, where the current flow is diffusive, the agreement between theory curves and measured data is no longer perfect. Thereby, that the best agreement occurs for the symmetric contact configuration in Fig. 9(a). This behavior suggests, that our model is inaccurate in the diffusive transport regime. Thereby, the deviations depend on the details of the contact configuration.
In conclusion, our four-terminal measurements confirm the main findings of our three-terminal measurements: (1) Away from the plateaus, the transport is not chiral. (2) For the plateau regions, the agreement between our four-terminal measurements and the chiral and coherent model supports chiral transport.
@@ III.4 Limits of our model in the diffusive regime
The deviations between the prediction of Eq. (II.2) and our measurements in the diffusive regime point to an incomplete model for $I_{xx}$. The problem is not related with the different contact resistances and offset voltages of the current amplifiers, because they are taken into account in Eq. (II.2). Instead, it is related with the mesoscopic dynamics of the electrons inside the Hall bar. In Sec. II.2 we determined the distribution of $I_{xx}$ into various contacts based on the contact resistances and the Hall resistance in series, while treating the Hall bar itself as a current node. To do this, we assumed a homogeneous current density distribution across the Hall bar. However, in particular near the intersections between the Hall bar and its contacts the distributions of the electric field and the current density become locally inhomogeneous. We conjecture, that such local inhomogeneities near the intersections cause the observed deviations between measured data and model curves in Fig. 9. Note that the inhomogeneities affect the distribution of $I_{xx}$ into various contacts, while $I_{xy}$ is nevertheless bent into the contacts due to the charge accumulated along the edge of the Hall bar. A more realistic model would require a suitable numerical calculation, e.g., based on a path integral formalism with realistic boundary conditions.
III.5 Experimental confirmation of non-chiral currents between plateaus
@@
Figure 10: Comparison of the respective largest measured current contributions flowing into a grounded contact for three different configurations of four-terminal measurements. The dotted horizontal lines depict the predicted current ratios for filling factors $1\leq\nu\leq 4$. The current ratios $I_{k}\,/\,I_{1}$ marked with a * are slightly scaled to account for the slightly different contact resistances, such that all three theory curves according to Eq. (II.2) with $\mu=\mu_{0}$ are identical and follow the edge between shaded and white backgrounds. Sketches: Arrows indicate the displayed current branches; colored numbers indicate current carrying and gray numbers floated contacts.
we plot for each of the three four-terminal measurements shown in Fig. 9 only the branch $I_{k}/I_{1}$ corresponding to the largest current. In a chiral model, this is the current flowing into the first grounded contact after emission from contact 1. We slightly scaled the data (based on the values provided in Table 1) to compensate for small variations in the contact resistances. After scaling, the three curves are identical in the classical regime (B<0.5B<0.5\,T) as well as on the plateaus of the QHE. The line linking shaded and white backgrounds corresponds to the prediction of Eq. (II.2) for constant $\mu=\mu_{0}$ [equal to the dashed black lines in Fig. 9(a)].
Compared to the model prediction, between the plateaus the current is reduced for all three geometries, which above we already interpreted in terms of a finite contribution of the non-chiral $I_{xx}$. Because the various contact resistances of our sample are almost identical and the number of grounded contacts is identical, for a purely chiral current flow the branches of the largest current $I_{k}(B)/I_{1}$ should be independent of the details of the contacts configuration. While this is the case on the quantized Hall plateaus, away from the plateuas $I_{k}(B)/I_{1}$ varies substantially between configurations. This can only be understood, if we allow for non-chiral currents between the quantized Hall plateaus.
This conclusion experts additional support by a closer look at the amounts of current reduction, based on its systematic dependence on the relative distances between contact 1 and the grounded contacts on both sample edges in a way that contradicts chiral current flow. Namely, the current supression into the contact with the largest current is the greater the closer another grounded contact is to the biased contact 1.
@@ III.6 Formulation of the proof of chiral current flow for quantized Hall resistance
So far, we have proven non-chiral current flow between the quantized plateaus, while for the quantized plateaus we have merely stated that our experimental results are consistent with chiral current flow. However, our analysis allows a stronger statement, namely that the agreement between the chiral model and our measurements is a prove of chiral current flow, if we accept that Eq. (3) describes chiral current flow.
The proof is based on the coherent and chiral model introduced in Sec. II.1. Based on the Landauer-Büttiker formalism it describes the current division into $n$ contacts in terms of the mutual conductance coefficients $G_{kj}$ between the contacts. In this model, chiral current flow corresponds to the specific values of conductance coefficients given by Eq. (3). A non-chiral behavior would correspond to different values, including additional non-zero conductance coefficients in Eq. (3). These would inevitably result in different plateau values of the predicted current ratios.
In all our measurements including three or four terminals shown in Fig. 7, Fig. 9 and Fig. 10, we instead see a perfect agreement between our chiral model prediction and the measured plateau values of the current ratios. This is, in particular, evident when comparing the horizontal dotted lines in Fig. 7(b) and Fig. 10, which indicate the predictions of Eq. (4) for the corresponding integer filling factors, with the corresponding measured plateau values.
@@ IV Summary and Outlook
We have performed multiterminal current measurements of a Hall bar in the regime of the QHE. We find that the current flow is chiral in the magnetic field regions of the quantized Hall resistance plateaus but becomes non-chiral in between the plateaus. Our result is based on the direct comparison of the measured current distribution into various contacts (three or four terminals) with two limiting models, each providing a unique physical meaning. The classical Drude model serves us as the basis to describe non-chiral diffusive transport, while our second model is based on the generic chiral and coherent transmission problem. The comparison of different contact configurations for more than three terminals allows for a direct experimental proof of non-chiral currents, which confirm our model based results. Our work contributes to the understanding of the current flow inside mesoscopic devices in the regime of the QHE. It thereby confirms the assumption of chiral current flow frequently used for the case of quantized plateaus of the Hall resistance.